OpenStudy (anonymous):

the sum of the multiples of 7 between 0 and 1000?

OpenStudy (amistre64):

\[\frac{1-7^n}{1-7}\]

OpenStudy (amistre64):

or am i thinking awry

OpenStudy (amistre64):

oh im thinking awry lol

OpenStudy (amistre64):

7n < 1000 n = 1000/7 = 142.xxxx 7*142 = 994

OpenStudy (amistre64):

142(7+994) ----------- should be close 2

OpenStudy (amistre64):

71071 if i did mine right

OpenStudy (anonymous):

this is an arithmetic series a1 = 7 and last term is 994 number of terms = 142 sum = 142/2 * [7 + 994] = 71 * 1001

OpenStudy (anonymous):

yep amis - thats it

OpenStudy (amistre64):

yay!!

OpenStudy (anonymous):

no more awry!!!!

OpenStudy (anonymous):

i think you thought geometric at first , right?

OpenStudy (amistre64):

yep